2.

Che 2 Problem 6-124 Each of the independent situ each year. The lessee is aware of the lessors implicit interest rate. ations below describes a finance lease in which annual lease payments are payable at the beginning of eBook Situation Print Lease term Lessors desired rate of return Lessees incremental borrowing rate Amount of lease 1eyrs 16% 2eyrs 12% 18% References 12% $600,000 $408,00e For convenience, here are some table values Periods; int. rateP PV, ordinary annuity PV, annuity 6.7590 10 periods, 10% le periods, 12% 28 periods, 1e 20 periods, 12% 6.1446 5.6582 8.5136 7.4694 6.3282 8.3658 Required: For each situation determine the amount recorded as a liability by the lessee at the beginning of the lease Amount of Liability Situation 1 Situation 2



7.

media%2F563%2F5638f3a8-0cc3-4c02-92e1-e3
Answer #1
A B C D E F G H I J
2
3 Calculation of Annual Payment:
4
5 Annual payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows:
6
7 Given the following data:
8 Loan Amount $225,000,000
9 Duration 12
10 Interest rate 8.00%
11
12 Annual Payment can be calculated as below:
13 RATE (Annual interst rate): 8.00%
14 NPER (No of Months): 12
15 PV (Loan Amount): -$225,000,000
16 FV 0
17 TYPE 0
18 Annual Payment $29,856,378.81 =PMT(D13,D14,D15,D16,D17)
19
20 Hence Annual Payment is $29,856,378.81
21

Formula sheet

A B C D E F G H I J
2
3 Calculation of Annual Payment:
4
5 Annual payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows:
6
7 Given the following data:
8 Loan Amount 225000000
9 Duration 12
10 Interest rate 0.08
11
12 Annual Payment can be calculated as below:
13 RATE (Annual interst rate): =D10
14 NPER (No of Months): =D9
15 PV (Loan Amount): =-D8
16 FV 0
17 TYPE 0
18 Annual Payment =PMT(D13,D14,D15,D16,D17) =PMT(D13,D14,D15,D16,D17)
19
20 Hence Annual Payment is =D18
21